The remainder of n7 is 4. The remainder of m9 is 7. Which of the following is a potential value for n×m?

(A) 24

(B) 44

(C)22

(D) 126

(E) 55


1
Expert's answer
2021-08-24T08:24:04-0400

Let "n=7x+4, x=0,1,2,..." and let "m=9y+7,y=0,1,2,..." Then


"n\times m=(7x+4)(9y+7)=63xy+49x+36y+28"

"y=0," if "n\times m" is divided by "7" the remainder is "0."

"126\div7=18"

"m=9(0)+7=7, n=7(2)+4=18"

"y=1," if "n\times m" is divided by "7" the remainder is "1."

"22\div7=3, remainder \ 1"

"m=9(1)+7=16"

But "22" is not divisible by "16."


"y=2," if "n\times m" is divided by "7" the remainder is "2."

"44\div7=6, remainder \ 2"

"m=9(2)+7=25"

But "44" is not divisible by "25."



"y=3," if "n\times m" is divided by "7" the remainder is "3."

"24\div7=2, remainder \ 3"

"m=9(3)+7=34"

But "24" is not divisible by "34."


"y=4," if "n\times m" is divided by "7" the remainder is "4."

"44\div7=6, remainder \ 2"

"m=9(2)+7=25"

But "44" is not divisible by "25."



"55\div7=7, remainder \ 6"

"y=6," if "n\times m" is divided by "7" the remainder is "6."

"m=9(6)+7=61"

But "55" is not divisible by "61."


(D) 126